calculate volume showing uncertainty;

L= 3.125 cm + 0.025 mm

W= 3.125 cm + 0.025 mm

h= 1.205 cm +0.025 mm
Answer #1

Volume(without uncertainity), V = L*W*h = 3.125*3.125*1.205 = 11.767 cm3

When quantities are multiplied, their percentage errors are added

percent errors

in L = (0.0025/3.125)*100 = 0.08 %

in W = (0.0025/3.125)*100 = 0.08 %

in h = (0.0025/1.205)*100 = 0.207 %

total error in volume = 0.08 + 0.08 + 0.207 = 0.367%

hence error in value in Volume = 3.67*11.767/100 = 0.043 cm3

hence error with uncertainity is 11.767 cm3 + 0.43 cm3